OpenStudy (anonymous):

Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2. y = 4x − 28 y = 4x + 28 y = 1 over 4x − 2 y = 1 over 4x + 2

OpenStudy (blank ):

A. y=4x-28

OpenStudy (anonymous):

Thank you so much! Can you explain it?

OpenStudy (blank ):

y=4x+2_(8,4) Solve the second equation to setup the y=mx+b formula. y=mx+b Using the y=mx+b formula, m=4. m=4 Find the value of b using the formula for the equation of a line. y=mx+b Substitute the value of y into the equation. (4)=(4)*(8)+b Find the value of b. b=-28 Now that the values of m(slope) and b(y-intercept) are known, substitute them into y=mx+b to find the equation of the line. y=4x-28 Hope this helps. :)

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (anonymous):

THANKS U SAVED MY LIFE

OpenStudy (anonymous):

wingspan